# Solute Transport due to Periodic Loading in a Soft Porous Material

### v1

@@ III.5 Quantitative impacts of periodic loading on solute transport
We next consider the effects of the loading parameters $A$ and $T$. To help interpret these results, we provide a brief summary of the key findings of our companion study before considering solute transport and mixing.
Poromechanical response to periodic loading
@@
As noted above, the poromechanical response impacts solute transport through the motion of the fluid and through the changes in porosity. In figure 7 and figure 8, respectively, we show the impact of $T$ and $A$ on the fluid flux $q_{f}$ (driving advection) and on $|v_{f}-v_{s}|$ (driving dispersion). As should be expected, the fluid flux exhibits localisation for fast loading and asymmetry in loading and unloading for large amplitudes. In details, in figure  7, we fix $A$ to the baseline value and consider four values of $T$. For slow loading, $|v_{f}-v_{s}|\sim A/T$ (see equation (47) and Ref. Fiori et al. (2023)). For very large values of $T$ (e.g. figure 7d and h), the deformation is uniform and very slow, and $|v_{f}-v_{s}|$ is low, especially towards the right (lighter lines) where the solute is positioned. Diffusion dominates over dispersion, even with a large amplitude. As $T$ decreases from $10\pi$ to $0.1\pi$ (e.g., figure 7b,c,f,g), the entire domain is characterised by an increasingly higher $|v_{f}-v_{s}|$. As $T$ decreases further, however, the deformation is increasingly localised at the left boundary: figure 7a and e, show that both $q_{f}$ and $|v_{f}-v_{s}|$ are orders of magnitude larger at the left boundary (darkest curves) than near the right boundary (lightest curves). This localisation is highlighted in figure 7i, where we plot the maximum value of $|v_{f}-v_{s}|$ at $x=1-l$ against $T$: towards the right boundary, where the solute is located, the strength of dispersion increases and then decreases with $T$, exhibiting a maximum around $T=0.1\pi$. In figure 8, we show the impact of $A$ on the same quantities for a fixed period, $T=0.1\pi$. As $A$ increases, $q_{f}$ and $|v_{f}-v_{s}|$ progressively increase, with increasingly larger and lower magnitudes during loading and unloading, respectively. For fixed $T$, the strength of dispersion increases monotonically with $A$ (figure 8i).
@@
Figure 7: Evolution of (a) – (d) fluid flux $q_{f}$ and (e) – (h) $|v_{f}-v_{s}|$ at ten different values of $X=x-u_{s}(X,t)$ from 00 to $1$ (dark to light) during one cycle for four different values of $T$ (columns). We distinguish between the loading half of the cycle ($\dot{a}>0$; solid curves) and the unloading half of the cycle ($\dot{a}<0$; dashed curves). (i) Maximum in time of $|v_{f}-v_{s}|$ at $x=1-l$ as a function of $T$.
@@
Figure 8: Evolution of (a) – (d): fluid flux $q_{f}$ and (e) – (h) $|v_{f}-v_{s}|$ at ten different values of $X=x-u_{s}(X,t)$ from 00 to $1$ (dark to light) during one cycle for $T=0.1\pi$, and for four different values of $A$ (columns). We distinguish between the loading half of the cycle ($\dot{a}>0$; solid curves) and the unloading half of the cycle ($\dot{a}<0$; dashed curves). (i) Maximum in time of $|v_{f}-v_{s}|$ at $x=1-l$ as a function of $A$.
Solute transport for different loading amplitude and period
@@
Figure 9: (a) Travel distance $\delta_{\mathcal{D}}$ and (b) degree of mixing $\chi_{\mathcal{D}}$ as a function of $T$ and for a wide range of $A$ (evenly spaced from 0.004 to 0.2, increasing dark to light) after a loading time of $12\pi$ followed by a relaxation time of 1 (total time $112\pi+$ 1) for advection and molecular diffusion but no dispersion ($\alpha=0$). Note that the values of $T$ are selected to provide an integer number of loading cycles in a total time of $12\pi$, but the results are plotted as continuous curves for visual clarity. This constraint leads to periods ranging from $T=0.01\pi$ applied for 1200 cycles to $T=12\pi$ applied for 1 cycle. Dashed lines (darkest color) correspond to diffusion with no loading ($A=0$). Note that the minimum porosity in the domain occurs near the piston and decreases monotonically with increasing $A$ and with decreasing $T$. Each curve ends on the left at the value of $T$ for which the minimum porosity vanishes and the simulations fail.
@@
We showed in section III.3 how the three transport mechanisms act individually on the concentration profile. We now extend this analysis to examine the role of $A$ and $T$. In figure 9, we show the travel distance $\delta_{\mathcal{D}}$ and the degree of mixing $\chi_{\mathcal{D}}$ after a fixed total total loading time of $12\pi$ followed by a relaxation time of 1 (total time $112\pi+$ 1), for advection and molecular diffusion but no dispersion ($\alpha=0$). We include results over a wide range of $T$ — from very fast to slow loading — and $A$ — from small to large deformations.
@@
We illustrated in section III.3 that the contribution of advection is reversible over one loading cycle, independent of $A$ and $T$. However, as noted above, both the porosity field and the concentration gradients do depend on $A$ and $T$. Hence, molecular diffusion is expected to vary weakly with $A$ and $T$. The porosity $\phi_{f}$ is on average lower than the initial value $\phi_{f,0}$, because the loading has a non-zero mean — the material is on average compressed. As noted in Ref. Fiori et al. (2023), the overall average porosity $\langle\overline{\phi}_{f}\rangle$ over any integer number of cycles is given by:
$\langle\overline{\phi_{f}}\rangle=\frac{1}{mT}\int_{nT}^{(n+m)T}\langle{\phi_{f}}\rangle\,\mathrm{d}t=1-\frac{1-\phi_{f,0}}{\sqrt{1-A}},$ (37)
@@
where $\langle{\phi_{f}}\rangle$ indicates the spatial average. Note that $\langle\overline{\phi}_{f}\rangle$ decreases with $A$ but is independent of $T$, as is also true of $\delta_{\mathcal{D}}$ and $\chi_{\mathcal{D}}$ in figure 9, except for very fast loading. In figure 10, we fix $T=12\pi$ and plot the normalised values of $\delta_{\mathcal{D}}$, $\chi_{\mathcal{D}}$, and $\langle\overline{\phi}_{f}\rangle$ against $A$, demonstrating that all three quantities exhibit a qualitatively similar decay with $A$.
Figure 10: Normalised $\delta_{\mathcal{D}}$ (dashed red curve), $\chi_{\mathcal{D}}$ (dotted blue curve), and $\langle\overline{\phi}_{f}\rangle$ (solid black curve) against $A$ for $T=12\pi$.
@@
Figures 9 and 10 confirm that $\delta_{\mathcal{D}}$ and $\chi_{\mathcal{D}}$ decrease weakly but monotonically with $A$, and are essentially independent of $T$ for all but the smallest values of $T$. For those smallest values, both $\delta_{\mathcal{D}}$ and $\chi_{\mathcal{D}}$ increase with $T$ (see (37)). This effect is not related to $\langle\overline{\phi}_{f}\rangle$, which is independent of $T$. We explore this behaviour in figure 11 by plotting the evolution of $\delta_{\mathcal{D}}$ over time for a fixed amplitude $A=0.06$ and for several of the smallest values of $T$, from $T=0.015\pi$ to $T=0.8\pi$.
@@
Figure 11: Time evolution of $\delta_{\mathcal{D}}$ at the end of a long series of cycles for six smallest values of $T$ considered here, $T=0.015\pi$ to $0.8\pi$ (dark to light), with advection and molecular diffusion but no dispersion ($\alpha=0$). The inset focuses on the very last portion of the main plot. Black stars mark the end of the last cycle of periodic loading and the beginning of the relaxation phase (total time of 1), during which the material returns to its undeformed state. Note that the horizontal axis is on a log scale.
@@
For the slowest case ($T=0.8\pi$), $\delta_{\mathcal{D}}$ is minimum at the end of the loading time ($12\pi$, marked by a black star) and then increases due to diffusion during the relaxation time (a further time of 1). For smaller values of $T$, $\delta_{\mathcal{D}}$ instead decreases during the initial stages of relaxation, immediately after the end of loading. This decrease is due to the relaxation of the static far-field compression induced for very fast loading (e.g., $T\lesssim 0.1\pi$). Since the deformation is much faster than the material response, the domain is never fully relaxed during periodic loading and the right portion, in particular, is in a state of static compression that contributes to an additional stretch of the concentration gradients that hinders diffusion. Once the loading stops, this compressed material relaxes by drawing fluid in, thereby further retracting the solute profile to a degree that is stronger and longer as $T$ decreases, such that $\delta_{\mathcal{D}}$ decreases very weakly as $T$ decreases in very fast loading (figure 10).
@@
Figure 12: (a) Travel distance $\delta$ and (b) degree of mixing $\chi$ (right) as a function of $T$ and for a wide range of $A$ (increasing dark to light) after a loading time of $12\pi$ followed by a relaxation time of 1 (total time $112\pi+$ 1), as in figure 9, but now with all three transport mechanisms simultaneously active. The ranges of amplitudes and periods are the same as in figure 9. Portions shown in greyscale indicate simulations where the solute reaches the left boundary and begins to leave the domain.
@@
We next repeat the preceding analysis, but now including hydrodynamic dispersion (figure 12). Both $\delta$ and $\chi$ are greatly enhanced by dispersion relative to the results in figure 9 across much of the range of $T$. Recall that the strength of dispersion is expected to scale with $A/T$. For very slow loading ($T>12\pi$), the contribution of dispersion is negligible relative to that of diffusion and the values of $\delta$ and $\chi$ converge toward their values without dispersion. As $T$ decreases, the contribution of dispersion grows and increasingly dominates over diffusion, reaching a peak around $T\approx 0.1\pi$. For these parameters, the values of $\delta$ in the peak region are almost one order of magnitude larger than without dispersion (cf. figure 9). As $T$ further decreases, $\delta$ and $\chi$ instead begin to decrease and again approach their values without dispersion for very fast loading ($T<~{}0.01\pi$). Across the full range of $T$ where dispersion dominates, $\delta$ and $\chi$ increase with $A$. This trend can be linked to the impact of $T$ on the ratio of $|v_{f}-v_{s}|$ at $x=1-l$, shown in figure 7(i): for faster loading, the deformation is increasingly localised near the piston and the region occupied by the solutes is increasingly not engaged. We examine these observations in more detail in figure 13, by plotting the evolution of $\delta$ throughout the loading time for a fixed amplitude and for several periods between $T=0.03\pi$ and $T=0.8\pi$, thus spanning the peaks in figure 12. In all cases, $\delta$ exhibits oscillations with period $T$ on top of a roughly Fickian growth. As $T$ decreases, these oscillations decrease in magnitude as they increase in frequency, consistent with the deformation being increasingly localised at the left. The overall rate of spreading increases as $T$ decreases from $0.8\pi$ to $0.1\pi$, for which the increase in frequency leads to a net increase in dispersive flux despite the decrease in magnitude (see figure 7). As $T$ decreases further, the decrease in local magnitude begins to dominate the increase in frequency and the rate of spreading instead slows (see again figure 7).
@@
We next compare the numerical values of $\delta$ shown in figure 12 with a naive analytical estimate $\delta_{\mathrm{est}}$ that includes both dispersion and localisation.
@@
Figure 14: Estimated travel distance $\delta_{\mathrm{est}}$ as a function of $T$ and for a wide range of $A$ (increasing dark to light) after a loading time of $12\pi$ followed by a relaxation time of 1 (total time $112\pi+$ 1), as in figure 12, but here calculated via equation (39).
@@
Following Ref. Fiori et al. (2023), the analytical solution for the Darcy flux during very fast loading at very low amplitude ($T\ll 1$ and $A\ll 1$; from linear poroelasticity and analogous to Stokes’s classical “second problem”) is
$\phi_{f}(v_{f,\mathrm{vf}}-v_{s,\mathrm{vf}})=-\frac{\pi A}{T}\left[\exp\left(-x\sqrt{\frac{\pi}{T}}\right)\sin\left(\frac{2\pi{}t}{T}-x\sqrt{\frac{\pi}{T}}\right)\right],$ (38)
@@
where the subscript (“$\mathrm{vf}$”) stands for “very fast”. Equation (38) suggests that, in this regime, the amplitude of the Darcy flux decays exponentially with distance from the left boundary and with $T^{-1/2}$. Based on this solution, we reformulate equation (36) to include localisation:
@@
$\delta_{\mathrm{est}}=C_{1}\sqrt{4\left[\mathrm{Pe}^{-1}+C_{2}\alpha\frac{A}{T}\exp\left({-(1-l)\sqrt{\frac{\pi}{T}}}\right)\right]t}.$ (39)
@@
Figure 14 shows that $\delta_{\mathrm{est}}$ captures the qualitative trends observed for $\delta$ in figure 12, with the central peak that is however larger and slightly shifted compared to the one observed for $\delta$. We do not expect strong quantitative agreement because this estimate ignores the details of the spatial variation in the dispersive flux, which decreases linearly from left to right even for slow loading, and also the oscillatory nature of the flow. Hence, figures 1214 highlight two competing mechanisms: progressively faster loading enhances dispersion by promoting large dispersive fluxes in general, but also progressively localises the flow and deformation near the left boundary, suppressing the flux (and thus dispersion) in the interior. The competition between these two behaviours is the origin of the local maximum in $\delta$ and $\chi$ with $T$.
@@
In summary, this analysis reveals the link between transport fluxes and loading parameters and, consequently, with the poromechanics of periodic loading. We identify two principal regions of the $A-T$ domain shown in Figure 12:
@@
  • $T\gtrsim 0.1\pi$: the loading is slow, and the strength of dispersion is roughly proportional to $A/T$. Thus, faster and larger deformations progressively increase the strength of dispersion relative to diffusion. In particular, as $T$ decreases $|v_{f}-v_{s}|$ reaches higher peaks, promoting hydrodynamic dispersion, which is the dominant mechanism for this region.

### v2

@@ III.5 Quantitative impacts of periodic loading on solute transport
We next consider the effects of the loading parameters $A$ and $T$. To help interpret these results, we first consider the poromechanical response.
III.5.1 Poromechanical response to periodic loading
@@
As noted above, the poromechanical response impacts solute transport through the motion of the fluid and through the changes in porosity. In figure 7 and figure 8, respectively, we show the impact of $T$ and $A$ on the fluid flux $q_{f}$ (driving advection) and on the relative velocity $|v_{f}-v_{s}|$ (driving dispersion). As should be expected, the fluid flux exhibits localisation for fast loading and asymmetry in loading and unloading for large amplitudes.
@@
Figure 7: Evolution of (a)–(d) fluid flux $q_{f}$ and (e)–(h) $|v_{f}-v_{s}|$ at ten different values of $X=x-u_{s}(X,t)$ from 00 to $1$ (dark to light) during one cycle for four different values of $T$ (columns). We distinguish between the loading half of the cycle ($\dot{a}>0$; solid curves) and the unloading half of the cycle ($\dot{a}<0$; dashed curves). (i) Normalised time-average of $|v_{f}-v_{s}|$ as a function of $X$ for the same four values of $T$ (increasing dark to light). (j) Maximum in time of $|v_{f}-v_{s}|$ at $x=1-l$ as a function of $T/\pi$ for $A=0.05$. The dashed black curve shows the slow-loading prediction of $\pi{}Al/(\phi_{f,0}T)$ and the dotted black curve shows the very-fast-loading prediction of $[\pi{}A/(\phi_{f,0}T)]\exp[-(1-l)\sqrt{\pi/T}]$ (see equation (42)).
In figure 7, we fix $A$ to the baseline value and consider four values of $T$. For slow loading, $|v_{f}-v_{s}|\sim 2A(1-x)/T$ (see Ref. Fiori et al. (2023)). For very large values of $T$ (e.g., figure 7d and h), the deformation is uniform and very slow, and $|v_{f}-v_{s}|$ is low, especially toward the right (lighter shades) where the solute is positioned. Diffusion dominates over dispersion, even with a large amplitude. As $T$ decreases from $6\pi$ to $0.1\pi$ (e.g., figure 7b,c,f,g), $|v_{f}-v_{s}|$ increases throughout the domain. As $T$ decreases further, however, the deformation is increasingly localised near the left boundary: figure 7a and e show that both $q_{f}$ and $|v_{f}-v_{s}|$ are orders of magnitude larger near the left boundary (darkest curves) than near the right boundary (lightest curves). This localisation is highlighted in figure 7i, where we plot the time-averaged profile of $|v_{f}-v_{s}|$ for each $T$, and in figure 7j, where we plot the maximum value of $|v_{f}-v_{s}|$ at $x=1-l$ against $T/\pi$. Near the right boundary, where the solute is located, the relative velocity increases and then decreases with $T$, exhibiting a maximum around $T=0.1\pi$.
@@
Figure 8: Evolution of (a)–(d) fluid flux $q_{f}$ and (e)–(h) relative velocity $|v_{f}-v_{s}|$ at ten different values of $X=x-u_{s}(X,t)$ from 00 to $1$ (dark to light) during one cycle for $T=0.1\pi$, and for four different values of $A$ (columns). We distinguish between the loading half of the cycle ($\dot{a}>0$; solid curves) and the unloading half of the cycle ($\dot{a}<0$; dashed curves). (i) Normalised time-average of $|v_{f}-v_{s}|$ as function of $X$ for the same four values of $A$ (increasing dark to light). (j) Maximum in time of $|v_{f}-v_{s}|$ at $x=1-l$ as a function of $A$.
In figure 8, we show the impact of $A$ on the same quantities for a fixed period, $T=0.1\pi$. The magnitudes of $q_{f}$ and $|v_{f}-v_{s}|$ increase monotonically with $A$ (figure 8a–h) and the normalised time-averaged value of $|v_{f}-v_{s}|$ is relatively insensitive to $A$ (figure 8i), suggesting that $|v_{f}-v_{s}|$ is essentially proportional $A$. This suggestion is confirmed in figure 8j.
III.5.2 Solute transport for different loading amplitudes and periods
@@
Figure 9: (a) Travel distance $\delta_{\mathcal{D}}$ and (b) degree of mixing $\chi_{\mathcal{D}}$ as a function of $T$ and for a wide range of $A$ (evenly spaced from 0.004 to 0.2, increasing dark to light) after a loading time of $12\pi$ followed by a relaxation time of 1 (total time $112\pi+$ 1) for advection and molecular diffusion but no dispersion ($\alpha=0$). Note that the values of $T$ are selected to provide an integer number of loading cycles in a total time of $12\pi$, but the results are plotted as continuous curves for visual clarity. This constraint leads to periods ranging from $T=0.01\pi$ applied for 1200 cycles to $T=12\pi$ applied for 1 cycle. Dashed lines (darkest color) correspond to diffusion with no loading ($A=0$). Note that the minimum porosity in the domain occurs near the piston and decreases monotonically with increasing $A$ and with decreasing $T$. Each curve ends on the left at the value of $T$ for which the minimum porosity vanishes and the simulations fail (see section II.4).
@@
We showed in section III.3 how the three transport mechanisms act individually on the concentration profile. We now extend this analysis to examine the roles of $A$ and $T$. In figure 9, we show the travel distance $\delta_{\mathcal{D}}$ and the degree of mixing $\chi_{\mathcal{D}}$ after a fixed total loading time of $12\pi$ followed by a relaxation time of 1 (total time $112\pi+$ 1), for advection and molecular diffusion but no dispersion ($\alpha=0$). We include results over a wide range of $T$ — from very fast to slow loading — and $A$ — from small to large deformations.
@@
We illustrated in section III.3 that the contribution of advection is reversible over one loading cycle, independent of $A$ and $T$. However, as noted above, both the porosity field and the concentration gradients do depend on $A$ and $T$. Hence, molecular diffusion is expected to vary weakly with $A$ and $T$. The porosity $\phi_{f}$ is on average lower than the initial value $\phi_{f,0}$, because the loading has a non-zero mean — the material is on average compressed. As noted in Ref. Fiori et al. (2023), the overall average porosity $\overline{\langle{\phi_{f}}\rangle}$ over any integer number of cycles is given by
$\overline{\langle{\phi_{f}}\rangle}=\frac{1}{mT}\int_{nT}^{(n+m)T}\langle{\phi_{f}}\rangle\,\mathrm{d}t=1-\frac{1-\phi_{f,0}}{\sqrt{1-A}},$ (41)
@@
where $m$ is an arbitrary positive integer and $n$ is an arbitrary non-negative integer. Note that $\overline{\langle{\phi_{f}}\rangle}$ decreases with $A$ but is independent of $T$, as is also true of $\delta_{\mathcal{D}}$ and $\chi_{\mathcal{D}}$ in figure 9, except for very fast loading. For slow loading, it is straightforward to show that the deformation stretches concentration gradients by, on average, a factor of $\phi_{f,0}/\overline{\langle{\phi_{f}}\rangle}$.
In figure 10a, we fix $T=12\pi$ and plot $\delta_{\mathcal{D}}$, $\chi_{\mathcal{D}}$, and $\overline{\langle{\phi_{f}}\rangle}$ against $A$; all three quantities are normalised by their values at $A=0$, demonstrating that they exhibit a qualitatively similar decay with $A$. We then plot $\delta_{\mathcal{D}}/\delta_{\mathcal{D},A=0}$ against $A$ for several values of $\phi_{f,0}$ (figure 10b) and $\mathrm{Pe}^{-1}$ (figure 10c). Panel (b) shows that increasing $\phi_{f,0}$ strongly mitigates the decay with $A$, while panel (c) shows that increasing $\mathrm{Pe}^{-1}$ has a similar but much weaker effect.
With no dispersion ($\alpha=0$), the ratio of $\delta_{\mathrm{est,SL}}$ to its value for $A=0$ is precisely $f\left(\overline{\langle{\phi_{f}}\rangle}/\phi_{f,0},\mathrm{Pe}^{-1}\right)$. Thus, the curves in figures 10b,c correspond to profiles of $f$ against $A$ for different values of $\phi_{f,0}$ and $\mathrm{Pe}^{-1}$. As indicated by the functional dependence of $f$, we hypothesise that $f$ depends on both $A$ and $\phi_{f,0}$ exclusively through the quantity $\overline{\langle{\phi_{f}}\rangle}/\phi_{f,0}$, which, as noted above, measures the (inverse of the) average stretching of concentration gradients due to the deformation (see section III.3). As indicated in Equation (40), we find that $f$ is linear in $\overline{\langle{\phi_{f}}\rangle}/\phi_{f,0}$ to a very close approximation; Figure 10b demonstrates excellent agreement between this expression for $f$ and our numerical results at fixed $\mathrm{Pe}^{-1}$.
The dependence of $f$ on $\mathrm{Pe}^{-1}$ is weaker. As indicated in Equation (40) and illustrated in Figure 10c, we find that a logarithmic variation with $\mathrm{Pe}^{-1}$ provides reasonable qualitative agreement between our expression for $f$ and our numerical results, with good quantitative agreement for larger values of $\mathrm{Pe}^{-1}$. Note that our expression for $f$ has just one fitting parameter, $C_{3}$, the value of which is strongly constrained by the functional form of $f$ and by the clear linear trend with $\overline{\langle{\phi_{f}}\rangle}/\phi_{f,0}$; of course, a better fit is possible with a different functional form and more fitting parameters.
Although the precise functional form chosen here for $f$ is ultimately ad hoc, it is clear that stronger stretching of the concentration profile (i.e., increasing $A$ or $\phi_{f,0}$, decreasing $\overline{\langle{\phi_{f}}\rangle}/\phi_{f,0}$) will more strongly hinder diffusion (decreasing $f$). The role of $\mathrm{Pe}^{-1}$ is more difficult to rationalise. We find that stretching of concentration gradients more strongly hinders diffusion at smaller values of $\mathrm{Pe}^{-1}$, corresponding to weaker diffusion. At smaller $\mathrm{Pe}^{-1}$, the initially steep concentration gradient will decay more slowly and it may be the case that diffusion is more sensitive to the stretching of these steeper gradients than shallower ones.
Figure 10: (a) Variation of $\delta_{\mathcal{D}}$ (red), $\chi_{\mathcal{D}}$ (blue), and $\overline{\langle{\phi_{f}}\rangle}$ (black) with $A$ for $T=12\pi$, and where all three quantities are normalised by their values at $A=0$. We also plot the variation of $\delta_{\mathcal{D}}$ with $A$, again normalised by its value at $A=0$, for (b) five values of $\phi_{f,0}$$\in[0.5,0.95]$ (light to dark) and (c) five values of $\mathrm{Pe}^{-1}$$\in[3\times 10^{-6},3\times 10^{-4}]$ (light to dark). Black dashed lines are the empirical function $f(\overline{\langle{\phi_{f}}\rangle}/\phi_{f,0},\mathrm{Pe}^{-1})$ from $\delta_{\mathrm{est,SL}}$ (see eqs. (39) and (40)).
@@
Figures 9 and 10 confirm that $\delta_{\mathcal{D}}$ and $\chi_{\mathcal{D}}$ decrease weakly but monotonically with $A$, and are essentially independent of $T$ for all but the smallest values of $T$. For those smallest values, both $\delta_{\mathcal{D}}$ and $\chi_{\mathcal{D}}$ increase with $T$. This effect is not related to $\overline{\langle{\phi_{f}}\rangle}$, which is independent of $T$ (see eq. (41)). We explore this behaviour in figure 11 by plotting the evolution of $\delta_{\mathcal{D}}$ over time for a fixed amplitude $A=0.06$ and for several values of $T$, from $T=0.015\pi$ to $T=0.8\pi$.
@@
Figure 11: Time evolution of $\delta_{\mathcal{D}}$ at the end of a long series of cycles for six of the smallest values of $T$ considered here, $T=0.015\pi$ to $0.8\pi$ (dark to light) with advection and molecular diffusion but no dispersion ($\alpha=0$). The inset focuses on the very last portion of the main plot. Black stars mark the end of the last cycle of periodic loading and the beginning of the relaxation phase (total time of 1), during which the material returns to its undeformed state. Note that the horizontal axis is on a log scale.
@@
For the slowest case ($T=0.8\pi$), $\delta_{\mathcal{D}}$ is minimum at the end of the loading time ($12\pi$, marked by a black star) and then increases due to diffusion during the relaxation time (a further time of 1). For smaller values of $T$, $\delta_{\mathcal{D}}$ instead decreases during the initial stages of relaxation, immediately after the end of loading. This decrease is due to the relaxation of the static far-field compression induced for very fast loading (e.g., $T\lesssim 0.1\pi$). Since the deformation is much faster than the material response, the domain is never fully relaxed during periodic loading and the right portion, in particular, is in a state of static compression that contributes to an additional stretch of the concentration gradients that hinders diffusion. Once the loading stops, this compressed material relaxes by drawing fluid in, thereby further retracting the solute profile to a degree that is stronger and longer as $T$ decreases, such that $\delta_{\mathcal{D}}$ decreases as $T$ decreases in very fast loading (figures 9 and 11).
@@
Figure 12: (a) Travel distance $\delta$ and (b) degree of mixing $\chi$ as functions of $T$ and for a wide range of $A$ (increasing dark to light) after a loading time of $12\pi$ followed by a relaxation time of 1 (total time $112\pi+$ 1), as in figure 9, but now with all three transport mechanisms simultaneously active. The ranges of amplitudes and periods are the same as in figure 9. Portions shown in greyscale indicate simulations where the solute reaches the left boundary and begins to leave the domain.
@@
We next repeat the preceding analysis, but now including hydrodynamic dispersion (figure 12). Both $\delta$ and $\chi$ are greatly enhanced by dispersion relative to the results in figure 9 across much of the range of $T$. Recall that the strength of dispersion is expected to scale with $A/T$. For very slow loading ($T\gtrsim{}12\pi$), the contribution of dispersion is negligible relative to that of diffusion and the values of $\delta$ and $\chi$ converge toward their values without dispersion. As $T$ decreases, the contribution of dispersion grows and increasingly dominates over diffusion, reaching a peak around $T\approx 0.1\pi$. For these parameters, the values of $\delta$ in the peak region are almost one order of magnitude larger than without dispersion (cf. figure 9). Across the full range of $T$ where dispersion dominates, $\delta$ and $\chi$ increase with $A$, as should be expected from figure 8(i): The relative velocity in the interior increases monotonically with $A$. As $T$ further decreases, $\delta$ and $\chi$ instead begin to decrease and for very fast loading ($T\lesssim{}0.01\pi$) approach their values without dispersion. This trend can be linked to the impact of $T$ on the ratio of $|v_{f}-v_{s}|$ at $x=1-l$, shown in figure 7(i): for faster loading, the deformation is increasingly localised near the piston and the region occupied by the solutes is increasingly not engaged. We examine these observations in more detail in figure 13, by plotting the evolution of $\delta$ throughout the loading time for a fixed amplitude and for several periods between $T=0.03\pi$ and $T=0.8\pi$, thus spanning the peak in figure 12. In all cases, $\delta$ exhibits oscillations with period $T$ on top of a roughly Fickian growth. As $T$ decreases, these oscillations decrease in magnitude as they increase in frequency, consistent with the deformation being increasingly localised at the left. The overall rate of spreading increases as $T$ decreases from $0.8\pi$ to $0.1\pi$, for which the increase in frequency leads to a net increase in dispersive flux despite the decrease in magnitude (see figure 7). As $T$ decreases further, the decrease in local magnitude begins to dominate the increase in frequency and the rate of spreading instead slows (see again figure 7).
@@
We next compare the numerical values of $\delta$ shown in figure 12 with a revised estimate $\delta_{\mathrm{est}}$ that accounts for diffusion, dispersion, and localisation.
@@
Figure 14: Estimated travel distance $\delta_{\mathrm{est}}$ as a function of $T$ and for a wide range of $A$ (increasing dark to light) after a loading time of $12\pi$ followed by a relaxation time of 1 (total time $112\pi+$ 1), as in figure 12, but here calculated via equation (43). The dashed black line corresponds to the maxima for $\delta$ from figure 12, for comparison.
@@
An analytical expression for the Darcy flux during very fast loading at very low amplitude ($T\ll 1$ and $A\ll 1$; from linear poroelasticity and analogous to Stokes’s classical “second problem”) can be derived from the results of Ref. Fiori et al. (2023) by calculating the solid velocity as $v_{s,\mathrm{vf}}\approx{}\mathrm{d}u_{s,\mathrm{vf}}/\mathrm{d}t$, where the solid displacement $u_{s,\mathrm{vf}}$ is given in their equation (3.12) and the subscript (“$\mathrm{vf}$”) stands for “very fast”. Equation (22) (also their equation 2.20) implies that the Darcy flux is then $\phi_{f,\mathrm{vf}}(v_{f,\mathrm{vf}}-v_{s,\mathrm{vf}})=-v_{s,\mathrm{vf}}$, such that
$\begin{split}\phi_{f,\mathrm{vf}}(v_{f,\mathrm{vf}}-v_{s,\mathrm{vf}})=-&\frac{\pi A}{T}\left[\exp\left(-x\sqrt{\frac{\pi}{T}}\right)\sin\left(\frac{2\pi{}t}{T}-x\sqrt{\frac{\pi}{T}}\right)\right].\end{split}$ (42)
@@
Equation (42) suggests that, in this regime, the amplitude of the Darcy flux decays exponentially with distance from the left boundary and with $T^{-1/2}$. Based on this solution, we reformulate equation (39) to include this exponential localisation in the magnitude of the dispersive flux:
@@
$\delta_{\mathrm{est}}=C_{1}f\left(\frac{\overline{\langle{\phi_{f}}\rangle}}{\phi_{f,0}},\mathrm{Pe}^{-1}\right)\sqrt{4\left[\mathrm{Pe}^{-1}+C_{2}\alpha\frac{A}{T}\exp\left({-(1-l)\sqrt{\frac{\pi}{T}}}\right)\right]t},$ (43)
@@
where $f\left(\frac{\overline{\langle{\phi_{f}}\rangle}}{\phi_{f,0}},\mathrm{Pe}^{-1}\right)$ is as defined in equation (40). Figure 14 shows that $\delta_{\mathrm{est}}$ captures the qualitative trends observed for $\delta$ in figure 12 (for which we report the maxima for a direct comparison), with the central peak that is however lower and slightly shifted compared to the one observed for $\delta$. We do not expect strong quantitative agreement because this estimate ignores the details of the spatial variation in the dispersive flux, which decreases linearly from left to right even for slow loading, as well as the oscillatory nature of the flow, and, most importantly, is based on an analytical solution that is only valid for small deformations and in the very-fast-loading region of the parameter space.
Hence, figures 1214 highlight two competing mechanisms: progressively faster loading enhances dispersion by promoting large dispersive fluxes in general, but also progressively localises the flow and deformation near the left boundary, suppressing the flux (and thus dispersion) in the interior. The competition between these two behaviours is the origin of the local maximum in $\delta$ and $\chi$ with $T$ and is qualitatively captured by the estimated travel distance $\delta_{\mathrm{est}}$.
@@
In summary, this analysis reveals the link between transport fluxes and loading parameters and, consequently, the poromechanics of periodic loading. Through the travel distance $\delta_{\mathrm{est}}$, we are able to capture the three main transport regimes, corresponding to analogous regions of the $A-T$ domain shown in Figure 12:
@@
  • $T\gtrsim 0.3\pi$: the loading is slow and the strength of dispersion is roughly proportional to $A/T$. Thus, faster and larger deformations progressively increase the strength of dispersion relative to diffusion. In particular, $|v_{f}-v_{s}|$ reaches higher peaks as $T$ decreases, promoting hydrodynamic dispersion, which is the dominant mechanism for this region.
  • $0.1\pi\lesssim T\lesssim 0.3\pi$: the loading is fast, with an optimal balance between magnitude and depth of the fluid fluxes; dispersion is at its peak.